# Shared Hardships Strengthen Bonds: Negative Shocks, Embeddedness and Employee Retention

### v1

@@ Abstract
Unexpected events—“shocks”—are the motive force in explaining changes in embeddedness and retention within the unfolding model of labor turnover. Substantial research effort has examined strategies for insulating valued employees from adverse shocks. However, this paper provides empirical evidence that unambiguously negative shocks can increase employee retention when underlying firm and employee incentives with respect to these shocks are aligned. Using survival analysis on a unique data set of 466,236 communication records and 45,873 employment spells from 21 trucking companies, we show how equipment-related shocks tend to increase the duration of employment. Equipment shocks also generate paradoxically positive sentiments that demonstrate an increase in employees’ affective commitment to the firm. Our results highlight the important moderating role aligned incentives have in how shocks ultimately translate into retention. Shared hardships strengthen bonds in employment as in other areas.
@@ 1 Introduction
In comparison to the national average voluntary separation rate of 2.2%, trucking industry turnover is extreme. [^1] December 2023; Latest figures: https://www.bls.gov/news.release/pdf/jolts.pdf These high rates of turnover provide variation that makes the trucking industry a particularly useful testbed for understanding successful worker retention strategies more generally. There are two important features of trucking, in particular, that make worker retention challenging. First, the job is highly-routinized across firms. Employees develop and depend little on firm-specific human capital so that their skills are highly substitutable between firms. In this environment, employers have little ability to offer firms-specific incentives to retain workers (Kryscynski et al., 2021). Second, the job is carried out remotely. While remote work, or work-from-anywhere policies have become important post-Covid, providing valuable amenities to workers and productivity gains to firms (Choudhury et al., 2021), this trend has a notable downside: it reduces employee embeddedness within firms because it makes it more difficult to build meaningful relational links with co-workers(Yang et al., 2022). These features constrain the strategies managers within the industry have for retention.
@@
With truck drivers less tied to their job through relational linkages, they are more sensitive to job “shocks”–random and often unexpected events that cause employees to re-examine their current employment relationship (Lee & Mitchell, 1994; Mitchell et al., 2001; Lee et al., 2004). Additionally, deprived of relational options for embedding employees within an organizations, managers may be tempted to treat or prevent the shock, rather than addressing the underlying job-fit characteristics, which the shock highlights. For example, recent studies have found shocks such as pay (Conroy et al., 2022) or schedule(Bergman et al., 2023) variance increase turnover. A naive policy would be to defend against these shocks, such as by reducing pay frequency so that workers experience fewer of these negative events. However, while this is sometimes the only option for managers, it is the equivalent of palliative care, failing to treat the fundamental job-fit mis-alignments the shock highlights.
@@
The main contribution of our paper is providing important empirical evidence that unambiguously deleterious events from either the firm or employee’s perspective, can result in increased retention and increased organizational commitment. We use data collected by a third-party human resources consultancy of weekly employee interviews of truck drivers for 21 trucking companies. We pair these weekly telephone interviews with employment records to follow over 45,873 employment spells. We aggregate communications records to identify 14 classes of shocks commonly faced by truck drivers, from equipment failures, and pay complaints, to issues of too little time-off. Using hazard analysis, we show that shocks where underlying incentives are aligned decrease driver turnover. In our sentiment analysis, equipment failures are followed by increased affective organizational commitment by drivers (Meyer & Allen, 1991). Our analyses use fixed effects and flexible hazard functions to control for heterogeneity across carriers, years, months, retention programs and agents. We also consider a number of alternative specifications and re-classifications of the shocks to demonstrate the robustness of our results.
@@
To understand the interesting pattern, we borrow the unfolding model of labor turnover (Lee & Mitchell, 1994), where “shocks” cause employees to re-examine their current employment relationships. From the perspective of the employer, these shocks can be “positive” (reducing turnover), or “negative” (increasing turnover). Scare quotes are necessary because the ultimate effect of shocks is moderated by the extent to which an employee is embedded within a firm, that is, how the shock itself highlights existing job fit, relational linkages, and sacrifices (Lee & Mitchell, 1994; Mitchell et al., 2001; Lee et al., 2004). Harmful events with respect to other firm or employee objectives can increase retention. For example, an equipment failure is bad news, both for the firm and the driver. Drivers are paid piece-rate, so that inoperative equipment reduces their ability to earn. The firm, as residual claimant in this context, has similar incentives towards maintaining equipment at a high-level. When equipment fails, it highlights an important area of mutual interest alignment between the driver and the motor carrier organization, paradoxically increasing retention.
@@
Our analysis yields a number of new and practical results. First, our extensive communications records allow us to assess the importance of a comprehensive set of work-related shocks on worker retention simultaneously. By far the most common, and the most strongly related to retention are equipment-related shocks (32%), and pay-related shocks (18%). Whereas shocks where interests are aligned, such as those related to equipment, tend to reduce turnover, pay-related shocks result in increased turnover precisely because interests are not aligned. We also contribute by performing a sentiment analysis allowing us to view the shocks from two distinct theoretical lenses (Philip & Medina-Craven, 2022): embeddedness theory (Mitchell et al., 2001), and organizational commitment(Meyer & Allen, 1991). Shocks that highlight interest alignment have positive associated sentiments, demonstrating increased commitment.
@@
The remainder of the paper is organized as follows. Section 2 reviews selected job turnover research. Formal hypotheses are developed in section 3. Section 4 discusses the dataset and empirical model. Section 5 presents results. Section 6 concludes the paper, summarizing our results, discussing research implications, and suggesting areas for future research.
@@ 2 Literature Review
Worker retention and turnover have been studied extensively in management(Holtom et al., 2008), sociology (Granovetter, 2017), psychology (Hom et al., 2017) and economics (Rogerson et al., 2005). Worker turnover is costly, incurring not only direct costs in worker replacement, but can also be accompanied by productivity losses (Ton & Huckman, 2008), reduced performance (Li et al., 2022; Hancock et al., 2013), product reliability concerns (Moon et al., 2022) and knock-on effects through contagion (Porter & Rigby, 2021). Understanding and reducing voluntary employee separation (hereafter, “job turnover”), is therefore a continuing concern. Our research contributes to two streams of research: employee turnover (Hom et al., 2017) and, more specifically, to understanding turnover within the trucking industry (Conroy et al., 2022).
2.1 Job Turnover
Several meta-analyses synthesize turnover predictors, including (Park & Shaw, 2013; Zimmerman & Darnold, 2009). Predictors can be associated with job, training, managerial, individual, and macroeconomic characteristics. Well studied job-related turnover determinants include role clarity, job routinization, promotion opportunities, performance and pay. The job characteristics, when combined with managerial aspects (satisfaction with leadership, influence in decision-making, team cohesion) and work-environment characteristics (stress, communication, distributive justice), determine key employee attitudes, which are strongly predictive of voluntary turnover, including overall job satisfaction, and organizational commitment (Allen et al., 2010).
Nonetheless, there are a number of empirical challenges that have been difficult to overcome. Much of the turnover literature relies on exit surveys, which must then deal with issues of recall(Hom, 2011) and selection bias (Pecker & Fine, 2015), and has been shown to be less useful than attitude survey of current employees (Fottler et al., 1995). In particular, reports from former employees can be biased by post-decision rationalizations or overstated or made-up justifications (Griffeth & Hom, 2001; Westaby, 2005). Moreover, causal empirical assessment requires defining a counterfactual control group (Morgan & Winship, 2007), which is impossible in exit surveys where all employees are in the ”treated” group of exiting employees.
@@ 2.1 Job Turnover
2.2 Trucking
@@
Turnover is a particular concern in the supply chain, especially in trucking, where extreme turnover rates generate extensive industry  [^2] https://www.trucking.org/news-insights/truth-about-trucking-turnover and academic interest (Miller et al., 2021).
@@
Studies have examined trucking turnover determinants that are common to other industries, for example, worker demographics, education and union status (Beilock & Capelle, 1990; Shaw et al., 1998), pay (LeMay et al., 1993; Phares & Balthrop, 2022; Conroy et al., 2022), tenure (Burks et al., 2006), economic climate (Prockl et al., 2017; Burks & Monaco, 2019; Miller et al., 2020), advancement opportunity (Min & Lambert, 2002), and fairness (Cantor et al., 2011). The logistics literature has also identified factors more particular to trucking, such as home time (Keller, 2002; Johnson et al., 2009, 2011; Burks et al., 2023) attitude towards dispatcher (Richard et al., 1995; Keller & Ozment, 1999; Morrow et al., 2005), lane balance (Sersland & Nataraajan, 2015) safety climate (Williams et al., 2011; Swartz et al., 2017) and emotional stress and loneliness (Kemp et al., 2013; Williams et al., 2017).
The self-stated purpose of much of this literature is to offer managerial insight, yet identifying antecedents to turnover is only the first step in designing programs to reduce turnover. Our study is the first to examine an explicit retention program, and is able to leverage a comparatively rich dataset on drivers’ self-reported concerns to uncover paradoxical correlations that highlighting the importance of interest alignment and job-fit in trucking.
3 Theory and Hypotheses
@@
We use the unfolding model of labor turnover (Lee et al., 2004, 1996) as our primary theoretical scaffolding for hypothesis development. In this framework, labor turnover proceeds along four possible pathways, which for density of exposition, we present out of numeric order. It can occur because of evolving job dissatisfaction by the worker (path 4) or be precipitated by shocks–unexpected events that cause employees to re-evaluate their current employment relationships (paths 1-3). Path 2 is a “push” driven path without a pre-planned response and without a specific job alternative in mind. An example might be being assigned to a new sales region when a worker hasn’t had this experience previously. The shock triggers a re-evaluation of current employment relationship which can result in exit. Path 1 is a “script-driven” pathway, where an event occurs that triggers a planned and immediate exit. Examples include surpassing a savings target or being assigned to a new sales territory (when a worker has prior experience with this event) could trigger a worker to leave employment. Path 3 encompasses “pull” shocks where a specific job alternative enters consideration, as when an employee becomes aware of another job possibility. For our purposes, the usefulness of the unfolding model is not in making precise distinction between shock pathways 1-3, but in understanding what different types of exogenous shocks reveal about retaining employees.
@@
Human resource managers can focus on two basic strategies for retaining employees: preventing or responding to shocks, or hardening employees against negative shocks. The latter strategy–the reason employees stay–is the complement of turnover, job embeddedness (Mitchell et al., 2001; Lee et al., 2014). Job embeddedness can be internal to the firm, or external, as when related to family, community, or other considerations. Dimensions of embeddedness applicable both internally and externally include links, fit and sacrifice.
@@
As a class, truck drivers are much less embedded within a job than workers in other occupations because external embeddedness considerations rarely fix workers to a particular job. For example, a truck driver does not necessarily need to move locations to take a job with a different company, even if that company is located far away. Truckers are also less internally embedded. Drivers work remotely, making them less beholden to relational links that are important in other occupations. Additionally, skills are highly transferable across firms so that there is comparatively little sacrifice involved in changing jobs. Truck driver embeddedness relies more narrowly on considerations of job fit, resulting in more sensitivity to shocks.
Truckers experience a variety of shocks in the course of their duties which may trigger reassessment of current employment including traffic congestion, equipment failures, detention during loading and unloading, variation in pay, and so on. The effect of the shock, however is moderated by embeddedness, and, in particular, what the shock reveals about job-fit. The ultimate effect of a shock, with respect to retention, is predictable based on interest alignment between firm and employee.
@@
There are shocks in trucking where drivers and employer interests are aligned. Equipment maintenance is one example. Long-haul truck drivers live out of their vehicles so that keeping vehicles in good condition contributes substantially to driver welfare. Drivers are liable for any maintenance problems that result in potential safety hazards. Additionally, because drivers are often paid piece-rate by mileage, any equipment failures that prevent vehicle operation cuts into the drivers’ pay. Drivers therefore have the incentive to immediately address and report equipment-related problems.
@@
Correspondingly, employers have the incentive to immediately respond and take substantive action to fix equipment-related problems. First, equipment-related problems that prevent a vehicle from operating hurt firms through reduced operational efficiency and potential late-delivery penalties from shippers. Second, severe equipment problems may result in safety violations, negatively affecting a firm’s safety record and thus their ability to do business.
We hypothesize, then, that equipment-related shocks underscore employee-employer interest alignment, which reduces turnover. Stated formally, we then have H1:
Hypothesis 1.
Shocks which highlight job-fit through interest alignment increase embeddedness, therefore reducing turnover.
Other examples of shocks where interests are aligned are in resolving operational inefficiencies, such a driver detention, inefficient routing, and so on.

### v2

@@ Abstract
Jarring events inspiring reflection, known as “shocks” in the literature, are the motive force in explaining changes in employee embeddedness and retention within the unfolding model of labor turnover. Substantial research effort has examined strategies for insulating valued employees from adverse shocks. However, this paper provides empirical evidence that unambiguously negative shocks can increase employee retention when underlying firm and employee incentives with respect to these shocks are aligned. Using survival analysis on a unique data set of 466,236 communication records and 45,873 employment spells from 21 trucking companies, we show how equipment-related shocks tend to increase the duration of employment. Equipment shocks also generate paradoxically positive sentiments that demonstrate an increase in employees’ affective commitment to the firm. Our results highlight the important moderating role aligned incentives have in how shocks ultimately translate into retention. Shared hardships strengthen bonds in employment as in other areas.
@@ 1 Introduction
In comparison to the national average voluntary separation rate of 2.2%, trucking industry turnover is extreme. [^1] December 2023; Latest figures: https://www.bls.gov/news.release/pdf/jolts.pdf These high rates of turnover provide variation that makes the trucking industry a particularly useful testbed for understanding successful worker retention strategies more generally. There are two important features of trucking, in particular, that make worker retention challenging. First, the job is highly-routinized across firms. Employees develop and depend little on firm-specific human capital so that their skills are highly substitutable between firms. In this environment, employers have little ability to offer firms-specific incentives to retain workers (Kryscynski et al., 2021). Second, the job is carried out remotely. While remote work, or work-from-anywhere policies have become important post-Covid, providing valuable amenities to workers and productivity gains to firms (Choudhury et al., 2021), this trend has a notable downside: it reduces employee embeddedness within firms because it makes it more difficult to build meaningful relational links with co-workers (Yang et al., 2022). These features limit the strategies managers within the industry have for retention.
@@
With truck drivers less tied to their job through relational linkages, they are more sensitive to the random and often unexpected events that cause employees to re-examine their current employment relationship. These events are known in the literature as “shocks” (Lee & Mitchell, 1994; Mitchell, Holtom, Lee, Sablynski & Erez, 2001; Lee et al., 2004). Additionally, deprived of relational options for embedding employees within an organizations, managers may be tempted to treat or prevent the shock, rather than addressing the underlying job-fit characteristics, which the shock highlights. For example, recent studies have found shocks such as pay (Conroy et al., 2022) or schedule (Bergman et al., 2023) variance increase turnover. A naive policy would be to defend against these shocks, such as by reducing pay frequency so that workers experience fewer of these negative events. However, while this is sometimes the only option for managers, it is the equivalent of palliative care, failing to treat the fundamental job-fit mis-alignments the shock highlights.
@@
The primary contribution of our paper is providing important empirical evidence that unambiguously deleterious events from either the firm or employee’s perspective, can result in increased retention and increased organizational commitment. We use data collected by a third-party human resources consultancy of weekly employee interviews of truck drivers for 21 trucking companies. We pair these weekly telephone interviews with employment records to follow over 45,873 employment spells. We aggregate communications records to identify 14 classes of shocks commonly faced by truck drivers, from equipment failures, and pay complaints, to issues of too little time-off. Through statistical analysis of employment durations, we show that shocks where firm and employee incentives are aligned decrease driver turnover. In our sentiment analysis, equipment failures are followed by increased affective organizational commitment by drivers (Meyer & Allen, 1991). Our empirical analyses use fixed effects and flexible model specifications to control for heterogeneity across carriers, years, months, retention programs and agents. We also consider a number of alternative specifications and re-classifications of the shocks to demonstrate the robustness of our results.
@@
To understand the interesting pattern, we borrow the unfolding model of labor turnover (Lee & Mitchell, 1994), where shocks cause employees to re-examine their current employment relationships. From the perspective of the employer, these shocks can be “positive” (reducing turnover), or “negative” (increasing turnover). Scare quotes are necessary because the ultimate effect of shocks is moderated by the extent to which an employee is embedded within a firm, that is, how the shock itself highlights existing job fit, relational linkages, and sacrifices (Lee & Mitchell, 1994; Mitchell, Holtom, Lee, Sablynski & Erez, 2001; Lee et al., 2004). Harmful events with respect to other firm or employee objectives can increase retention. For example, an equipment failure is bad news, both for the firm and the driver. Drivers are paid piece-rate, so that inoperative equipment reduces their ability to earn. The firm, as residual claimant in this context, has similar incentives towards maintaining equipment at a high-level. When equipment fails, it highlights an important area of mutual interest alignment between the driver and the motor carrier organization, paradoxically increasing retention.
@@
Our analysis yields a number of new and practical results. First, our extensive communications records allow us to assess the importance of a comprehensive set of work-related shocks on worker retention simultaneously. By far the most common are equipment-related shocks (33.7%), and pay-related shocks (18%). Whereas shocks where interests are aligned, such as those related to equipment, tend to reduce turnover, pay-related shocks result in increased turnover precisely because interests are not aligned. We also contribute by performing a sentiment analysis allowing us to view the shocks from two distinct theoretical lenses (Philip & Medina-Craven, 2022): embeddedness theory (Mitchell, Holtom, Lee, Sablynski & Erez, 2001), and organizational commitment (Meyer & Allen, 1991). Shocks that highlight interest alignment have positive associated sentiments, demonstrating increased commitment.
@@
The remainder of the paper is organized as follows. Section 2 reviews selected job turnover research. We develop our formal hypothesis in section 3. Section 4 discusses the dataset and empirical model. Section 5 presents results. Section 6 concludes the paper, summarizing our results, discussing research implications, and suggesting areas for future research.
@@ 2 Literature Review
Worker retention and turnover have been studied extensively in management (Holtom et al., 2008), sociology (Granovetter, 2017), psychology (Hom et al., 2017) and economics (Rogerson et al., 2005). Worker turnover is costly, incurring not only direct costs in worker replacement, but can also be accompanied by productivity losses (Ton & Huckman, 2008), reduced performance (Li et al., 2022; Hancock et al., 2013), product reliability concerns (Moon et al., 2022) and knock-on effects through contagion (Porter & Rigby, 2021). Understanding and reducing voluntary employee separation (hereafter, “employee turnover”), is therefore a continuing concern. Our research contributes to two streams of research: employee turnover (Hom et al., 2017) and, more specifically, to understanding turnover within the trucking industry (Conroy et al., 2022).
2.1 Employee Turnover
Several meta-analyses synthesize turnover predictors, including Park & Shaw (2013) and Zimmerman & Darnold (2009). Predictors can be associated with job, training, managerial, individual, and macroeconomic characteristics. Well studied job-related turnover determinants include role clarity, job routinization, promotion opportunities, performance and pay. The job characteristics, when combined with managerial aspects (satisfaction with leadership, influence in decision-making, team cohesion) and work-environment characteristics (stress, communication, distributive justice), determine key employee attitudes, which are strongly predictive of voluntary turnover, including overall job satisfaction, and organizational commitment (Allen et al., 2010).
Nonetheless, there are a number of empirical challenges that have been difficult to overcome. Much of the turnover literature relies on exit surveys, which must then deal with issues of recall (Hom, 2011) and selection bias (Pecker & Fine, 2015), and has been shown to be less useful than attitude surveys of current employees (Fottler et al., 1995). In particular, reports from former employees can be biased by post-decision rationalizations or overstated or made-up justifications (Griffeth & Hom, 2001; Westaby, 2005). Moreover, causal empirical assessment requires defining a counterfactual control group (Morgan & Winship, 2007), which is impossible in exit surveys where all employees are in the “treated” group of exiting employees.
@@ 2.1 Employee Turnover
2.2 Employee Turnover within the Trucking Industry
@@
Turnover is a particular concern in trucking, where extreme turnover rates generate extensive industry [^2] https://www.trucking.org/news-insights/truth-about-trucking-turnover and academic interest (Miller et al., 2021).
@@
Studies have examined trucking turnover determinants that are common to other industries, for example, worker demographics, education and union status (Beilock & Capelle, 1990; Shaw et al., 1998), pay (LeMay et al., 1993; Phares & Balthrop, 2022; Conroy et al., 2022), tenure (Burks et al., 2006), economic climate (Prockl et al., 2017; Burks & Monaco, 2019; Miller et al., 2020), advancement opportunity (Min & Lambert, 2002), and fairness (Cantor et al., 2011). The logistics literature has also identified factors more particular to trucking, such as home time (Keller, 2002; Johnson et al., 2009, 2011; Burks et al., 2023) attitude towards dispatcher (Richard et al., 1995; Keller & Ozment, 1999; Morrow et al., 2005), lane balance (Sersland & Nataraajan, 2015), safety climate (Williams et al., 2011; Swartz et al., 2017), and emotional stress and loneliness (Kemp et al., 2013; Williams et al., 2017).
Much of the previous retention literature relies on cross-sectional surveys of comparatively few respondents ($<600$), with the key dependent variable being intention to leave, e.g., Prockl et al. (2017); Swartz et al. (2017); Schulz et al. (2014). This has a few drawbacks. Stated preference (intention) can be poorly indicative of actual behaviors (Chandon et al., 2005), and causal assessment from cross-sectional data requires comparatively strong assumptions about the data generating process (Winship & Morgan, 1999). A separate stream of research focuses on macroeconomic aggregates and market conditions (Miller et al., 2021; Burks & Monaco, 2019; Phares & Balthrop, 2022; Burks et al., 2023), using individual and industry-level information to understand turnover. These studies do not have the same level of intra- and cross-firm variation and lack the granularity to identify the effect of specific shocks on retention.
This paper fills an important gap. Our dataset has significant within and cross-firm variation, unique in the literature. Additionally, its panel structure allow for us to control for unobserved time-invariant firm heterogeneity. Indeed, the data uncovers significant cross-firm variation in retention. Weekly call intervals and the rich characterization of driver complaints allows us to address previously identified retention correlates simultaneously, tracing out the evolving effect of shocks on retention. The outcome is that we are able to uncover paradoxical correlations that highlight the importance of interest alignment and job-fit within the trucking industry.
3 Theory and Hypothesis
3.1 The Unfolding Model of Labor Turnover and Job Embeddedness
We use the unfolding model of labor turnover (Lee et al., 2004, 1996) as our primary theoretical scaffolding for hypothesis development. This theory unifies two classes of explanations for voluntary employee turnover: external market driven reasons (“pull-related”), such as higher-wage employment options in other occupations (Phares & Balthrop, 2022) and internal psychological-push reasons related to the employee (“push-related”), such as too little home time (Williams et al., 2017). The unfolding model draws on image theory (Beach, 1990), in which decisions are made through a screening process as opposed to an exhaustive evaluation of options, as under expected utility theory. Agents, awash in information and stimuli, evaluate a subset of the stimuli, determining whether the events are consistent with their objectives, or require behavioral adjustments. The unfolding model posits that a worker’s decision to leave a job can be more readily explained by the arrival of new impetuses, triggering reflection and job re-evaluation, rather than by objective changes in the ecosystem of employment options. The stimulus must be sufficiently jarring to give the employee cause for reflection, an event referred to as a “shock.”
@@ 3.1 The Unfolding Model of Labor Turnover and Job Embeddedness
Labor turnover proceeds along four possible pathways in the unfolding model. Exit along pathways 1-3 is triggered by shocks. Path 1 is a “script-driven” automatic pathway, where an event occurs that triggers a planned and immediate exit. Examples include surpassing a savings target, or being assigned to a new sales territory when a worker has had prior experience with this event. Path 2 is a “push” driven path without a pre-planned response and without a specific job alternative in mind. An example might be being assigned to a new sales region when a worker has not had this experience previously. Path 3 encompasses “pull” shocks where a specific job alternative enters consideration, as when an employee becomes aware of another job possibility. Here it is important to emphasize that it is the awareness of the job alternative, not the job alternative itself, that explains retention. In path 4, a worker decides to leave because of evolving job dissatisfaction. Path 1 exits are the most immediate. Path 2-3 exits requires time for reflection and decision, occurring over a few days to a few weeks. Path 4 exits unfold over the longest horizon of months and years (Mitchell, Holtom & Lee, 2001). For our purposes, the usefulness of the unfolding model is in its dynamical treatment of shocks and the comprehensive event-centered organizing principles designed to provide tactical guidance to managers in day-to-day business operations.
@@
Human resource managers can focus on two basic strategies for retaining employees: preventing or responding to shocks, or hardening employees against negative shocks. The latter strategy–the reason employees stay–is the complement of turnover, job embeddedness (Mitchell, Holtom, Lee, Sablynski & Erez, 2001; Lee et al., 2014). Job embeddedness can be internal to the firm, or external, as when related to family, community, or other considerations. Dimensions of embeddedness applicable both internally and externally include links, fit and sacrifice.
  • Links are people-centered and include the relationships with colleagues, mentors, and management within the organization. Employees with many links within an organization stay longer. Links can also be external to the organization, such as family considerations which keep an employee at a job, or links to industry or trade groups, and non-work friendships.
  • Fit regards an employee’s perceived satisfaction within an organization including whether the employee is able to achieve personal goals, whether values and interests are aligned between the individual and organization, and whether the employee has the requisite skills, training and experience in their particular role. External to the organization, fit encompasses the degree to which the employee is comfortable outside of the organization, including the enjoyment of local amenities, local political and religious attitudes, and so on.
  • Sacrifice conceptualizes how easy an individual finds it to dissolve and rebuild links and fit. Sacrifice also embodies the switching cost between jobs. Other things equal, employees are more likely to remain in their current position if leaving involves significant loss. For example, employees only partially vested in the company retirement plan will find it more costly to leave than when they become fully vested.
Research has shown embeddedness both mediates and moderates the effect of shocks, with complex relationships with performance and organizational citizenship (Lee et al., 2014; Jiang et al., 2012). Shocks cause reflection on links, fit and sacrifice, changing embeddedness (mediation). When links and fit are especially strong, shocks have a much weaker effect on the continued employment decision than otherwise (moderation). Conversely, for weakly embedded employees, comparatively mundane events can trigger employee exit.
Job embeddedness is a latent construct whereas shocks are observable. To measure embeddedness, researchers have used measures solicited from surveys designed for the purpose (Mitchell & Lee, 2001; Crossley et al., 2007). The concept of embeddedness has been useful in a variety of job settings, with embeddedness firmly established as an important correlate of turnover in meta-analyses (Park & Min, 2020; Peltokorpi et al., 2015; Jiang et al., 2012). As with other archival research, we do not observe embeddedness directly. Instead, we use embeddedness theory as an organizing framework for understanding how context specific forces and shocks translate into retention through the unfolding model of labor turnover (Holtom & Inderrieden, 2006).
3.2 Hypothesis Development
@@
As a class, truck drivers are much less embedded within a job than workers in other occupations because external embeddedness considerations rarely fix workers to a particular job. For example, a truck driver does not need to move locations to take a job with a different company, even if that company is located far away. Truckers are also less internally embedded. Drivers work remotely, making them less beholden to relational links that are important in other occupations. Additionally, skills are highly transferable across firms so that there is comparatively little sacrifice involved in changing jobs. Truck driver embeddedness relies more narrowly on considerations of job fit, resulting in less job attachment and more sensitivity to shocks.
The types of shocks that are relevant to retention are specific to the particular circumstance of the job. Nurses are exposed to a different vector of shocks in the line of their duties than teachers or law enforcement officers. For trucking, we focus on 14 categories of shocks commonly faced by truckers, which have been highlighted by previous academic and industry research. These shocks are summarized in Table 1, including categories related to Home time, Waiting, Pay, Equipment, Safety/Ethics, Training, Operations, Disrespect, Company Policy, COVID, Personal and Health. Superficially, the shocks would seemingly negatively effect retention.
  • Drivers work remotely away from home and family. Routes and loadings are often planned in advance to minimize empty truck miles, leaving the driver committed to an itinerary. Shocks involving home issues, such as a missed family event, an emergency requiring the driver’s presence at home, or an unanticipated load pick-up that extends the driver’s time away cause the driver to reflect on their current employment relationship along paths 1 and 2 and 4 in the unfolding model. Previous research has demonstrated the importance of home time to driver retention (Taylor et al., 2010; Williams et al., 2011; Johnson et al., 2009, 2011; Burks et al., 2023).
  • Pay is the most obvious shock related to path 3 in the unfolding model, although it can also involve other paths. Pay-related issues are commonly cited within the retention literature (Allen et al., 2010). For tucking specifically, where drivers are paid piece-rate by mileage, pay issues can also manifest as scheduling and mileage issues that keep drivers from achieving earnings targets (Williams et al., 2011; Prockl et al., 2017; Conroy et al., 2022).
  • Wait time affects driver’s time away from home and their realized hourly wage rate. Generally, truckers are not compensated while they wait to be loaded or unloaded, or during mandatory rest intervals, increasing uncompensated hours (Belzer, 2000). Wait time also lowers the firm’s effective equipment and driver utilization rates, diminishing firm earnings. Exploiting electronic logging device data, Correll (2024) finds that wait time is highly predictive of turnover. We expect wait-time to be primarily push-driven on pathway 2, but may also operate pathways 1 and 4.
  • Equipment breakdowns and maintenance-related issues prevent safe operations and inhibit the driver’s ability to earn. Trucks also provide shelter and amenities to drivers on long-haul routes, contributing significantly to ensure driver comfort. Investing in equipment is a long-known and common strategy for recruiting and retaining drivers (Rakowski et al., 1989; Williams et al., 2011). We expect equipment shocks to operate through pathways 1,2 and 4. Importantly, safe and consistent operation of equipment is also in the firm’s interest.
  • Safety and ethical ethical concerns are also a common reason for voluntary turnover (Williams et al., 2011; Swartz et al., 2017). These concerns include violation of government or company rules, such as exceeding hours-of-service limits (Scott et al., 2021), unsafe loading or driving conditions, or equipment that violates FMCSA rules (Balthrop et al., 2023). Serious concerns about unsafe vehicle operation and unethical business practices likely operate strongly along path 1 in the unfolding model of labor turnover. Less serious concerns involving some amount of reflection affect retention through paths 2 and 4.
  • Training concerns include complaints related to training pay, training-related accommodations and travel, training content and duration. The effect of training itself on retention can be ambiguous depending on its portability, visibility and credibility (Dietz & Zwick, 2021), but training-related complaints increase push-related turnover (Consenza et al., 2012; Williams et al., 2011) likely along paths 1,2, and 4.
  • Team problems include personal problems with teammate (e.g., hygiene), scheduling issues related to teaming, team-related training, and issues related to team lifestyle (e.g., cramped conditions). A good team environment provides relational links that may keep a driver at a firm, but team-related problems do the opposite. We expect team issues to be primarily push-related, operating through pathways 1,2 and 4.
  • Operations concerns, including issues such as routing, fueling, loading/unloading practices, and issues related to trailer pick-up and drop-off (Johnson et al., 2009) can act as shocks causing re-consideration of employment along push pathways 1,2 and 4.
  • Disrespect includes issues related to confrontational dispatchers and driver’s treatment by firms’ and customers’ employees. Disrespect is a pervasive driver experience (Williams et al., 2017), and may contribute to push-related exit on paths 1,2 and 4.
  • Company policies related to company size, culture, customers, and rules requiring driver-facing cameras and sleep apnea machines can be a source of irritation to drivers resulting in exit along paths 1, 2, and 4.
  • Covid. The Covid-19 pandemic provided a strong impetus for exiting (Sperry et al., 2022). Trucking was considered a frontline critical industry. But travel to hotspots, loss of access to supporting infrastructure (rest-stops, fueling stations, restaurants) and exposure to potential infection provided strong reason for drivers to leave along path 1 of the unfolding model.
  • Personal issues include family issues aside from home-time concerns, retirement decisions, preference for being a leased-operator, and commute to work. We expect personal issues to result in push-driven exit along paths 1,2, and 4.
  • Health concerns unrelated to Covid are also an important source of retention issues. Trucking is a stressful profession associated with a variety of physical and mental health concerns (Williams et al., 2017; Aryal et al., 2023). Health issues provide a script driven reason to exit along path 1.
We emphasize that the listed shocks are all negative events from the driver’s perspective. The ultimate effect of the shock on retention, however is influenced by embeddedness, and, in particular, what the shock highlights with respect to job-fit. The effect of a shock on retention is predictable based on interest alignment between firm and employee.
@@
There are shocks in trucking where drivers and employer interests are aligned. Equipment maintenance is one example. Long-haul truck drivers live out of their vehicles so that keeping vehicles in good condition contributes substantially to driver welfare (Williams et al., 2011). Drivers are liable for any maintenance problems that result in potential safety hazards. Additionally, because drivers are often paid piece-rate by mileage (Ju & Belzer, 2022), any equipment failures that prevent vehicle operation cuts into the drivers’ pay . Drivers therefore have the incentive to immediately address and report equipment-related problems.
@@
Correspondingly, employers have the incentive to immediately respond and take substantive action to fix equipment-related problems. First, equipment-related problems that prevent a vehicle from operating hurt firms through reduced operational efficiency and potential late-delivery penalties from shippers. Second, severe equipment problems may result in safety violations and even legal liabilities, negatively affecting a firm’s safety record and thus their ability to do business (Miller et al., 2022).
We hypothesize, then, that equipment-related shocks underscore employee-employer interest alignment, highlighting the employees sense of of job-fit, reducing turnover. Other examples of shocks where interests are aligned are in resolving operational inefficiencies, such a driver detention, inefficient routing, and wait time. For drivers constrained by hours-of-service regulation, operational inefficiencies and wait time reduce the effective hours drivers have to cover miles (Scott et al., 2021) resulting in reduced earnings. Again, the motor carrier firm’s incentives with respect to wait time and operation are aligned, because these inefficiencies reduce the productivity of both drivers and their equipment, reducing the firm’s earnings.
Interest alignment can also result in substantive leader-member exchange leaving the employee further embedded into the motor carrier (Lee et al., 2014). When equipment, operations or wait time shocks are addressed immediately and effectively by management, such a by the dispatch of maintenance for faulty equipment, or the adjustment of itineraries to minimize driver wait time at loading docks, the effect is to leave the employee feeling that they are an important part of business operations, and that their concerns are received by management and taken seriously. Other things equal, these type of shocks make it more likely for employees to stay.